OpenStudy (anonymous):

simplify the expression 27x^-3y^-7/36x^-3y

OpenStudy (campbell_st):

well its simply \[\frac{27}{36} \times \frac{x^{-3}}{x^{-3}} \times \frac{y^{-7}}{y^1}\] so simplify the numbers at the start then use the index law for division is subtract the powers \[\frac{x^a}{x^b} = x^{a -b}\] and you'll get the answer... remember the index law for a power of zero... \[a^0 = 1\] you'll need it for this question.

OpenStudy (anonymous):